# Contrastive Learning for Task-Independent SpeechLLM-Pretraining

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in translation and language understanding tasks (OpenAI et al., 2024; Dubey et al., 2024). However, their reliance on written text limits their application to real-world scenarios, where communication often occurs through speech. Extending LLMs to spoken language could enable them to reason about and process spoken content effectively.
@@
Pretraining the SpeechLLM before finetuning on task-specific data can help overcome this challenge and reduce the need for large downstream datasets. A straightforward strategy is to use ASR data, training on the ASR task itself (Tang et al., 2024; Wang et al., 2024b). However, this might lead to overfitting to this pretraining task.
@@
Scaling up contrastive pretraining to 1,400 hours yields additional performance gains, with our models surpassing specialized models and current SpeechLLMs. Even in simulated low-resource scenarios, where only 10% of downstream data is used for finetuning, our method consistently outperforms existing SpeechLLMs.
@@ 2 Related Works
Chu et al. (2024) adopt a comprehensive strategy by finetuning all components. Similarly, other works explore finetuning both the projector and the encoder, alongside applying Low-Rank Adaptation (LoRA; Hu et al. 2022) to the LLM (Wu et al., 2023; Gong et al., 2024). In contrast, some methods freeze the LLM entirely, focusing on training only the encoder and projector (Wang et al., 2024b; Held et al., 2024). Several other approaches focus on finetuning the projector while introducing LoRA layers to the LLM, leaving the speech encoder and LLM frozen (Peng et al., 2024; Tang et al., 2024). An even more lightweight approach freezes both the encoder and the LLM, finetuning only the projection layers (Wang et al., 2024a).
@@
While ‘alignment’ has various meanings (Hämmerl et al., 2024), in this work, we use it to refer to the process of creating meaningfully similar representations for text and speech. One approach to achieve this alignment is to directly finetune on the desired tasks without explicitly modelling alignment (Fang et al., 2024). An alternative is to use a downstream task to pretrain the SpeechLLM before finetuning on desired tasks, for example pretraining on ASR (Tang et al., 2024; Wang et al., 2024b) or ASR variants (Wang et al., 2024a).
@@ 3 Methods
We propose contrastive learning as a task-independent pretraining method to align speech and text representations, leveraging abundantly available ASR data. We compare with two ASR-based pretraining baselines: traditional ASR pretraining and next-word prediction with mixed text-speech spans, which we also introduce as a pretraining approach in this paper.
@@ 3.1 Contrastive Training
Contrastive training brings representations of matched speech-text pairs closer whilst pushing representations of mismatched pairs apart.
Harnessing both matched and mismatched pairs is often a very efficient way of using a dataset and allows one to leverage abundantly available ASR data.
@@
One way to compute $\text{sim}(\mathbf{s}_{i},\mathbf{t}_{i})$ is to use cosine similarity. Since the length of the speech sequences, $M$, is typically much larger than the length of the text sequences, $N$, we cannot directly compute cosine similarity on the entire sequence. Instead, we average the embeddings over the sequence length to obtain a single embedding for each (Ye et al., 2022). We denote this pretraining strategy with contr-cos.
@@
Using the Wasserstein distance in the loss is expensive, so we resort to an upper-bound approximation that can be efficiently computed using the Sinkhorn algorithm (Sinkhorn and Knopp, 1967).
We use the negative distance as the similarity measure for $\mathcal{L}_{\text{InfoNCE}}$ and denote this pretraining strategy with contr-wasser.
@@ 3.2 Other Pretraining Approaches
We evaluate our contrastive pretraining by comparing it to traditional ASR pretraining and a mixed text-speech NWP pretraining approach.
@@
Inspired by Nguyen et al. (2024), who add mixed NWP to their finetuning data mix for downstream tasks,
we propose mixed text and speech NWP as a pretraining strategy.
We mix text and speech embeddings across randomly selected spans in the sequence, maintaining coherence with word boundaries. A sample sequence might look as follows: $t_{0}^{0},t_{1}^{0},s_{2}^{1},s_{3}^{1},s_{4}^{2},s_{5}^{2},s_{6}^{2},t_{7}^{3}$, where superscripts represent word boundaries, and subscripts represent subwords within the text or speech embedding. Since speech embeddings typically differ in length from text embeddings, we compute the NWP loss exclusively on text tokens. Details on how we select the speech and text spans are provided in Appendix A.
@@ 4 Experiments
Our experiments follow a two-stage training strategy: first, we perform text-speech alignment pretraining using ASR data to adapt the model to handle speech. In the second stage, we finetune the model jointly on a mix of task-specific data. This approach leverages abundant ASR data to reduce reliance on large task-specific datasets, enabling effective adaptation even in low-resource settings.
@@ 4.1 Model Architecture
For contrastive learning, we compare speech and text embeddings without the use of system prompts. This creates a positional mismatch between contrastive pretraining and finetuning. To address this potential overfitting problem, we experiment with varying the absolute starting position in the positional embeddings through random sampling for each example during pretraining.
@@ 4.2 Training Parameters
In both pretraining and task-specific finetuning, we keep the LLM and speech encoder frozen, adjusting only the speech projector, which has 42.5 million parameters. Training is conducted on four NVIDIA A100-SXM4-40GB GPUs. ASR pretraining takes approximately 9h, contrastive pretraining with cosine similarity over all layers takes around 5h. The task-specific finetuning takes around 30h with 100% of the data. Detailed training parameters and model specifics are listed in Appendix B.
@@ 4.3 Data
We also analyse whether contrastive pretraining harms a model’s ability to capture paralinguistic features, such as speaking pace or noise. We use the mls-eng-speaker-descriptions dataset (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024), which involves reasoning about speaking rate, gender, background noise, and other features.
More information about the train and test sets can be found in Appendix C and details on prompts in Appendix D.
@@ 4.4 Evaluation
N/A Specialized 06.54 30.99 80.02 64.19 77.10 N/A N/A 100
N/A HuBERT + Llama 18.38 19.85 73.92 36.15 54.76 N/A N/A 0
100% no pretrain 12.09 28.84 79.94 62.39 74.44 1.37 1.26 79.96
10% no pretrain 23.78 19.87 69.72 31.35 41.86 1.38 1.36 -57.40
ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 39.93
contr-cos-emb 13.98 25.47 76.32 55.10 68.72 0.97 0.72 46.34
contr-wasser-emb 15.99 26.10 77.09 55.04 68.17 1.08 0.62 44.02
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 67.39
contr-wasser-all 12.92 29.07 80.52 64.04 77.26 1.07 0.64 85.02
@@
(a) contr-cos pretrained models
(b) contr-wasser pretrained models
Figure 2: Analysis of a contrastive pretraining performed on different layers. We report the normalized average over tasks as in Eq. 3 for contrastive pretraining and finetuning on 10% of task-specific data. The last point on the x-axis refers to contrastive pretraining using the loss sum over all layers.
@@ 5.1 Contrastive Pretraining
We apply contrastive pretraining to the embedding layer using either Wasserstein loss (contr-wasser-emb) or cosine similarity loss (contr-cos-emb). Models are then jointly finetuned on three downstream tasks using 10% of the data to simulate low-resource settings (Table 1).
Our contrastive pretraining models consistently outperform the cascaded Hubert+Llama and no-pretraining baselines across all metrics. In ST and SQA tasks, they also surpass standard ASR pretraining, by over 5% in SQA. However, ASR-pretrained models remain better for ASR tasks. Interestingly, while contrastive models exhibit lower test-set contrastive loss on the embedding layer, the losses for no-pretraining and ASR-pretraining are similar despite a large performance gap.
@@
Keeping with the low-resource setting, we investigate calculating the contrastive loss at different LLM layers. Figure 2 demonstrates that applying contrastive loss to deeper layers significantly increases downstream performance, with improvements of up to 20% for cosine similarity loss and 40% for Wasserstein loss.
Further analysis reveals that deeper layers benefit ST and SQA the most, while improvements for ASR are less consistent. We analyse both contrastive loss and task performance across layers in detail in Appendix E. To address positional mismatches between pretraining and finetuning (detailed in Section 4.1), we experiment with position shifts during pretraining. However, these do not consistently improve performance.
Contrastive Loss on All Layers.
Although Fig. 2 shows a clear advantage to applying the contrastive loss to deeper layers, the best layer to use is not consistent. Therefore, we also explore using a summed loss over all (multiples of 5) layers in Table 1. The methods contr-cos-all and contr-wasser-all achieve normalized task averages of $67.39$ and $85.02$, respectively, significantly outperforming the ASR-pretraining baseline and the contr-emb models. Contr-wasser-all, despite using only 10% of the finetuning data, even matches the performance of the specialized models on two out of three tasks and surpasses the no pretraining model finetuned on 100% of the task-specific data. This highlights the suitability of contrastive pretraining for scenarios with limited task-specific data.
@@ 5.2 Mixed-Text-Speech Input
no pretrain 23.78 69.72 41.86 -57.40
ASR pretrain 12.21 75.70 63.36 039.93
contr-cos-emb 13.98 76.32 68.72 046.34
   + mixed 14.26 76.61 68.40 046.61
contr-wasser-emb 15.99 77.09 68.17 044.02
   + mixed 14.70 75.05 61.48 026.53
contr-cos-all 13.06 78.29 73.90 067.39
   + mixed 17.63 73.83 55.12 002.14
mixed-nwp 15.27 76.05 67.68 0 39.67
@@
We adapt our approach to use mixed speech-text sequences (contr-*-*+mixed). Using the contr-cos loss on the embedding layer, our model matches the performance of standard contrastive pretraining, outperforming both no pretraining and ASR pretraining (Table 2). However, with the contr-wasser loss results degrade significantly. Consequently, we continue with contr-cos-all+mixed to test the loss across all layers. This model also significantly underperforms its non-mixed counterpart. Further ablations of contrastive pretraining with mixed speech-text inputs do not yield additional improvements (detailed results, including BLEU and EM scores, are provided in Appendix F).
Mixed NWP.
We also propose using a next-word prediction (NWP) objective with mixed speech-text inputs (mixed-nwp) for pretraining, inspired by Nguyen et al. (2024). While this approach surpasses the no pretraining and ASR pretraining baselines (Table 2, last row), it remains inferior to contrastive pretraining.
5.3 Combining losses
@@
ASR pretrain 12.21 75.70 63.36 39.93
contr-cos-emb 13.98 76.32 68.72 46.34
   + asr loss 13.18 76.89 69.23 52.47
contr-wasser-emb 15.99 77.09 68.17 44.02
   + asr loss 13.19 77.4 70.79 57.54
contr-cos-all 13.06 78.29 73.90 67.39
   + asr loss 11.68 78.94 76.40 78.58
contr-wasser-all 12.92 80.52 77.26 85.02
   + asr loss 11.23 80.26 77.57 88.80
@@
N/A Specialized 06.54 30.99 80.02 64.1 77.1 N/A N/A 100
N/A BLSP-lslm-7b 44.51 28.70 78.68 5.60 21.82 N/A N/A -96.72
N/A Qwen2-Audio-7b 12.03 21.57 74.81 27.79 46.75 N/A N/A 010.79
10% ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 039.93
   + giga 12.16 27.44 78.38 62.23 73.40 1.37 1.31 069.72
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 067.39
   + giga 10.94 29.95 81.21 65.58 79.32 1.06 0.71 097.43
contr-cos-all + asr 11.68 27.61 78.94 64.16 76.40 1.08 0.91 078.58
   + giga 11.12 29.90 81.29 72.43 82.63 0.98 0.65 102.30
100% ASR pretrain 10.28 30.31 80.98 70.06 81.01 1.37 1.22 100.53
   + giga 11.48 30.46 81.03 72.41 82.73 1.37 1.23 100.02
contr-cos-all 12.56 30.69 81.48 71.93 82.70 1.25 1.02 099.36
   + giga 9.31 31.36 81.98 75.28 84.74 1.10 0.83 114.29
contr-cos-all + asr 10.04 30.82 81.56 72.57 82.72 1.19 1.03 106.94
   + giga 10.03 31.54 81.99 76.25 85.01 1.10 0.79 112.76
We now explore the potential of combining the previous pretraining objectives.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in language understanding and translation tasks (OpenAI et al., 2024; Dubey et al., 2024). However, their reliance on written text limits their application in real-world scenarios, where communication often occurs in the form of speech. Extending LLMs to spoken language could enable them to reason about and process spoken content effectively.
@@
Pretraining SpeechLLMs before finetuning on task-specific data can help overcome this challenge, reducing the need for large downstream datasets. A straightforward strategy is to use ASR data, training on the ASR task itself (Tang et al., 2024; Wang et al., 2024b). However, this might lead to overfitting to this pretraining task (Tang et al., 2024).
@@
Scaling up contrastive pretraining to 1,400 hours yields further performance gains, with our models surpassing specialized models and current SpeechLLMs. Even in simulated low-resource scenarios, using only 10% of the finetuning data, our method consistently outperforms existing SpeechLLMs.
@@ 2 Related Works
Chu et al. (2024) adopt a comprehensive strategy by finetuning all components. Similarly, other works explore finetuning both the projector and the encoder, alongside applying Low-Rank Adaptation (LoRA; Hu et al. 2022) to the LLM (Wu et al., 2023; Gong et al., 2024; Fathullah et al., 2023). In contrast, some methods freeze the LLM entirely, focusing on training only the encoder and projector (Wang et al., 2024b; Held et al., 2024; Fathullah et al., 2024). Several other approaches focus on finetuning the projector while introducing LoRA layers to the LLM, leaving the speech encoder and LLM frozen (Peng et al., 2024; Tang et al., 2024). An even more lightweight approach freezes both the encoder and the LLM, finetuning only the projection layers (Wang et al., 2024a).
@@
While ‘alignment’ has various meanings (Hämmerl et al., 2024), in this work, we use it to refer to the process of creating meaningfully similar representations for text and speech. One approach to achieve this alignment is to directly finetune on the desired tasks without explicitly modelling alignment (Fang et al., 2024; Fathullah et al., 2023). An alternative is to use a downstream task to pretrain the SpeechLLM before finetuning on desired tasks, for example pretraining on ASR (Tang et al., 2024; Wang et al., 2024b) or ASR variants (Wang et al., 2024a).
@@ 3 Methods
We propose contrastive learning as a task-independent pretraining method for aligning speech and text representations, leveraging abundantly available ASR data. We compare with two ASR-based pretraining baselines: traditional ASR pretraining and next-word prediction with mixed text-speech spans, which we also introduce as a pretraining approach in this paper.
@@ 3.1 Contrastive Training
Contrastive training brings representations of matched speech-text pairs closer whilst pushing representations of mismatched pairs apart. Harnessing both matched and mismatched pairs is often a very efficient way of using a dataset and allows one to leverage abundantly available ASR data.
@@
One way to compute $\text{sim}(\mathbf{s}_{i},\mathbf{t}_{i})$ is to use cosine similarity. Since the length of the speech sequences, $M$, is typically much larger than the length of the text sequences, $N$, we cannot directly compute cosine similarity on the entire sequence. Instead, we average the embeddings over the sequence length to obtain a single embedding for each (Ye et al., 2022). We denote this pretraining strategy by contr-cos.
@@
Using the Wasserstein distance in the loss is expensive, so we resort to an upper-bound approximation that can be computed efficiently using the Sinkhorn algorithm (Sinkhorn and Knopp, 1967). Note, however, that even with this approximation, using Wasserstein distance is more computationally expensive ( $O(n^{2}/epsilon^{2})$ (Pham et al., 2020)) than using cosine similarity ($O(n)$).
We use the negative distance as the similarity measure for $\mathcal{L}_{\text{InfoNCE}}$ and denote this pretraining strategy by contr-wasser.
@@ 3.2 Other Pretraining Approaches
To evaluate our contrastive pretraining approach we compare it with traditional ASR pretraining and a mixed text-speech NWP pretraining approach.
@@
Inspired by Nguyen et al. (2024), who add mixed NWP to their finetuning data mix for downstream tasks, we propose mixed text and speech NWP as a pretraining strategy. We mix text and speech embeddings across randomly selected spans in the sequence, maintaining coherence with word boundaries. A sample sequence might look as follows: $t_{0}^{0},t_{1}^{0},s_{2}^{1},s_{3}^{1},s_{4}^{2},s_{5}^{2},s_{6}^{2},t_{7}^{3}$, where superscripts represent word boundaries, and subscripts represent subwords within the text or speech embedding. Since speech embeddings typically differ in length from text embeddings, we compute the NWP loss exclusively on text tokens. Details on how we select the speech and text spans are provided in Appendix E.
@@ 4 Experiments
Our experiments follow a two-stage training strategy. First, we perform text-speech alignment pretraining using ASR data to adapt the model to handle speech. In the second stage, we finetune the model jointly on a mix of task-specific data. This approach leverages abundant ASR data to reduce reliance on large task-specific datasets, enabling effective adaptation even in low-resource settings.
@@ 4.1 Model Architecture
For contrastive learning, we compare speech and text embeddings without the use of system prompts. This creates a positional mismatch between contrastive pretraining and finetuning. To address this potential overfitting problem, we experiment with varying the absolute starting position in the positional embeddings. However, this does not consistently improve model performance (Appendix D).
@@ 4.2 Training Parameters
In both pretraining and task-specific finetuning, we keep the LLM and speech encoder frozen, adjusting only the speech projector, which has 42.5 million parameters. Training is conducted on four NVIDIA A100-SXM4-40GB GPUs. ASR pretraining takes approximately 9h, contrastive pretraining with cosine similarity over all layers takes around 5h. The task-specific finetuning takes around 30h with 100% of the data. Detailed training parameters and model specifics are listed in Appendix A.
@@ 4.3 Data
While there is overlap between pretraining and finetuning data for the ASR and ST tasks, we consider this a realistic scenario: In practice, when task-specific data is limited, leveraging all available data—even if it appears in both stages—is often necessary. This is particularly relevant for speech translation, where transcripts are abundant but parallel translations are much scarcer.
@@
We also analyse whether contrastive pretraining harms a model’s ability to capture paralinguistic features, such as speaking pace, gender or noise. We use two datasets: the mls-eng-speaker-descriptions dataset (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024), which involves reasoning about speaking rate, gender, background noise, and other features, and the MuST-SHE (Bentivogli et al., 2020) dataset, where gender information is needed to correctly translate the segment.
More information about the train and test sets can be found in Appendix B and details on prompts in Appendix C.
@@ 4.4 Evaluation
We also report the contrastive loss on the test set for the embedding layer in order to evaluate the effectiveness of aligning speech and text.
@@
N/A Specialized 06.54 30.99 80.02 64.19 77.10 N/A N/A -100
N/A HuBERT + Llama 18.38 19.85 73.92 36.15 54.76 N/A N/A -0
100% no pretrain 12.09 28.84 79.94 64.13 76.53 1.37 1.26 -83.08
10% no pretrain 23.78 19.87 69.72 34.72 46.07 1.38 1.36 -51.13
ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 -39.93
contr-cos-emb 13.98 25.47 76.32 54.86 68.49 0.97 0.72 -46.00
contr-wasser-emb 15.99 26.10 77.09 54.64 67.83 1.08 0.62 -43.52
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 -67.39
contr-wasser-all 12.92 29.07 80.52 64.06 77.22 1.07 0.64 -84.96
@@ 5.1 Contrastive Pretraining
We start by calculating the contrastive loss on the embedding layer and comparing the two loss functions: Wasserstein (contr-wasser-emb), and cosine similarity (contr-cos-emb). After pretraining, the models are finetuned on three downstream tasks (jointly) using a random 10% subset of the data, simulating a low-resource setting (Table 1).
Our contrastive pretraining models consistently outperform the cascaded Hubert+Llama and no-pretraining baselines across all metrics. In ST and SQA, they also surpass standard ASR pretraining, by over 5% in SQA. However, ASR-pretrained models remain better for ASR, which is unsurprising, given they are pretrained with more data on this task. Analysing the contrastive loss on the test set, we find that it is significantly lower after contrastive pretraining than after ASR pretraining or direct finetuning on the task, showing that indeed there is more alignment. Interestingly, while contrastive models exhibit lower test-set contrastive loss on the embedding layer, the losses for no-pretraining and ASR-pretraining are similar despite a large performance gap.
@@
Keeping with the low-resource setting, we investigate calculating the contrastive loss at different LLM layers. We find that applying contrastive loss to deeper layers significantly increases downstream performance, however, the best layer to use is not consistent. Therefore, we also explore using a summed loss over all (multiples of 5222Initial experiments found that using every 5th layer efficiently approximates using all layers (Table 11).) layers in Table 1. The methods contr-cos-all and contr-wasser-all achieve normalized task averages of $67.39$ and $84.96$, respectively, significantly outperforming the ASR-pretraining baseline and the contr-emb models. Contr-wasser-all, despite using only 10% of the finetuning data, even matches the performance of the specialized models on two out of three tasks and surpasses the no pretraining model finetuned on 100% of the task-specific data. This highlights the suitability of contrastive pretraining for scenarios with limited task-specific data. We analyse the impact of training on different layers in more detail in Appendix D, including the impact on individual task performance.
@@ 5.2 Mixed-Text-Speech Input
no pretrain 23.78 69.72 46.07 -51.13
ASR pretrain 12.21 75.70 63.36 -39.93
contr-cos-emb 13.98 76.32 68.49 -46.00
   + mixed 14.26 76.61 68.40 -46.61
contr-wasser-emb 15.99 77.09 67.83 -43.52
   + mixed 14.70 75.05 61.48 -26.53
contr-cos-all 13.06 78.29 73.90 -67.39
   + mixed 17.63 73.83 55.12 -02.14
mixed-nwp 15.27 76.05 67.68 -39.67
@@
We adapt our approach to use mixed speech-text sequences (contr-*-*+mixed). Using the contr-cos loss on the embedding layer, our model matches the performance of standard contrastive pretraining, outperforming both no pretraining and ASR pretraining (Table 2). However, with the contr-wasser loss results degrade significantly. Consequently, we continue with contr-cos-all+mixed to test the loss across all layers. This model also significantly underperforms its non-mixed counterpart.
@@
ASR pretrain 12.21 75.70 63.36 -39.93
contr-cos-emb 13.98 76.32 68.49 -46.00
   + asr loss 13.18 76.89 68.67 -51.64
contr-wasser-emb 15.99 77.09 67.83 -43.52
   + asr loss 13.19 77.40 70.44 -57.03
contr-cos-all 13.06 78.29 73.90 -67.39
   + asr loss 11.68 78.94 76.12 -78.17
contr-wasser-all 12.92 80.52 77.22 -84.96
   + asr loss 11.23 80.26 77.19 -88.24
@@
N/A Specialized 06.54 30.99 80.02 64.19 77.10 N/A N/A -100
N/A BLSP-lslm-7b 44.51 28.70 78.68 05.60 21.82 N/A N/A 0-96.72
N/A Qwen2-Audio-7b 12.03 21.57 74.81 27.79 46.75 N/A N/A -010.79
10% ASR pretrain 12.21 24.82 75.70 49.48 63.36 1.37 1.30 0-39.93
   + giga 12.16 27.44 78.38 62.23 73.17 1.37 1.31 -069.37
contr-cos-all 13.06 27.19 78.29 60.48 73.90 1.08 0.91 -067.39
   + giga 10.94 29.95 81.21 65.45 79.12 1.06 0.71 -097.12
contr-cos-all + asr 11.68 27.61 78.94 63.81 76.12 1.08 0.91 -078.17
   + giga 11.12 29.90 81.29 72.16 82.52 0.98 0.65 -102.15
100% ASR pretrain 10.28 30.31 80.98 69.87 80.92 1.37 1.22 -100.24
   + giga 11.48 30.46 81.03 72.25 82.56 1.37 1.23 -099.77
contr-cos-all 12.56 30.69 81.48 71.71 82.50 1.25 1.02 -099.06
   + giga 09.31 31.36 81.98 75.12 84.67 1.10 0.83 -114.18
contr-cos-all + asr 10.04 30.82 81.56 72.40 82.47 1.19 1.03 -106.57
   + giga 10.03 31.54 81.99 76.11 84.94 1.10 0.79 -112.66
Table 4: Comparison of pretraining only on Must-C data or the combination with the bigger Giga dataset (+giga). We compare our models against BLSP (Wang et al., 2024a) and the Qwen2-Audio-7B-Instruct model (Chu et al., 2024). Metrics with * contribute to the normalized average as in Eq. 3. Results for contr-wasser can be found in Appendix H and do not outperform contr-cos.
Mixed NWP.
We also propose using next-word prediction (NWP) with mixed speech-text inputs (mixed-nwp) for pretraining, inspired by Nguyen et al. (2024). While this approach surpasses the no pretraining and ASR pretraining baselines (Table 2, last row), it underperforms contrastive pretraining.
Further ablations of contrastive pretraining with mixed speech-text inputs and of mixed-nwp, including two different forced-alignment algorithms, do not yield additional improvements (detailed results, including all metrics, are provided in Appendix F).
5.3 Combining losses
Building on the previous sections, which analyse contrastive pretraining and mixed speech-text pretraining in comparison to traditional ASR pretraining, we now explore the potential of combining these approaches. We sum the individual losses, assigning equal weight to each.